# Extracting Essential and Disentangled Knowledge for Recommendation Enhancement

### v1

@@ 5.3.1. Impact of the Two Proposed Principles (RQ2)
To enforce the parametric knowledge base to memorize the essential and disentangle knowledge patterns within limited model parameters, we propose two constraints to regularize the knowledge compression process. In this section, we investigate the effectiveness of the two proposed constraints. We conduct experiments with backbones on which EDK achieves the best performance. Concretely, we conduct experiments with the DeepFM backbone for AD and conduct experiments on the DCN backbone for eleme. We remove the regularizations of the essential and disentangled principles to see the impacts of them. The results are displayed in Table 3.
Table 3. Impact of the Proposed Principles.
Datasets AD Eleme
AUC Logloss AUC Logloss
EDK 0.6378 0.1935 0.6563 0.0881
w/o Disentangled 0.6356 0.1939 0.6433 0.0886
w/o Essential 0.6347 0.1937 0.6520 0.0884
w/o both 0.6304 0.1940 0.6411 0.0888
@@
From the results, we can see that the proposed essential and disentangled constraints improve the quality of the memorized knowledge of the parametric knowledge base. Both the essential principle and the disentangled principle can solely contribute to the final performance, while the two complement each other and can collaborate for more qualified knowledge.
5.3.2. Impact of the Knowledge Extractor (RQ3)
@@
The patterns of the old data space are diverse and complex, being hard to be fully memorized due to the existence of spurious patterns and entanglement. To memorize the informative patterns within the data space, a knowledge extractor is designed to extract the essential and disentangled patterns within data points. When the number of knowledge pattern per data instance is set to 1, it reduces to memorizing a raw data point in the space. In this section, we conduct experiments on the number of knowledge patterns per data point to study the impact of the knowledge extractor. Similarly, we select DeepFM as the backbone for AD dataset and DCN as the backbone for Eleme dataset. We test on the range of $\left[5,10,15,20,25\right]$, the results are listed in Figure 4.
@@
Figure 4. Performances of EDK w.r.t different number of knowledge patterns per sample.
@@
5.4. Objectives of the Principles (RQ4)
In this section, we analyze whether the proposed two principles are achieved during the knowledge compression process.
5.4.1. Loss Curves of Different Components
@@
Figure 5. The learning curves of different losses on AD dataset.
The learning curves of different losses on AD dataset is illustrated in Figure 5. The blue line displays the loss curve for the disentangled constraint, which is $\sum_{i\neq j}I(s_{i};s_{j})$. The red line and black line together make up the loss for the essential constraint, which is $-\alpha\sum_{j}I(\mathbf{s_{j}};y)+I(X;\mathbf{c})$. Concretely, the black line illustrates the curve for $-\sum_{j}I(\mathbf{s_{j}};y)$ which promotes the memorization for task relevant information, and the red line illustrates the curve for $I(X;\mathbf{c})$, which compresses the representation and helps to filter the noisy information.
From the learning curves, we can see that all losses are learned well. While for the red one which corresponds to the vCLUB value, the MI upper bound estimation increases at the first epoch and decrease in the following training steps, since it takes some steps first to train the MI estimator well for correct estimation.
@@ 5.4.2. Visualizations of the Extracted Patterns
Then, we visualize the extracted patterns for more explicit analysis. Concretely, we analyze the pattern masks generated by the knowledge extractor and the corresponding representation encoded by the knowledge encoder.
Firstly, we study the masks generated by the trained knowledge extractor. Note that the entries of the masks are continuous. Here we regard an entry with value larger than $0.5$ as an existing feature and calculate the number of the non-zero entries of the generated masks. The extracted pattern scale distribution is illustrated in Figure 6.
@@
Figure 6. Pattern scales for AD and Eleme. (Entry with mask value$>0.5$ is regarded as 1.)
From the results, we could find that the scales of extracted patterns vary. For Eleme, the extracted patterns include distinct features, feature interactions, and denoised views of the original inputs. While for AD, the extracted patterns mainly consist of high-order feature interactions and denoised views of the inputs.
@@
In addition, we further study the representations of the raw data instance features and the extracted patterns. We use the T-SNE visualization to illustrate the distribution of the encodings of raw data instances and the extracted pattern vectors as displayed in Figure 7.Each point in the left part of the figure denotes a representation of a raw data instance, and each point in the right figure represents the an encoded pattern vector. From the results, we could see that there are apparently different manifolds existing in the extracted patterns representation space for both AD and Eleme datasets, while the raw data instances just appear in different clusterings. This validates the effectiveness of the proposed method on extracting the diverse and complex patterns from the old data.
@@ 6. Conclusion
In this paper, we design a parametric knowledge base EDK that compresses the massive old data into compact knowledge stored in parameters, which is regularized by the proposed essential and disentangled principles. The two principles promote robust and generalized knowledge memorized by the parametric knowledge base without increasing the size of it. The essential principle strikes a balance between compression and prediction, which filters out the noisy and irrelevant information of the old data and preserves the task-relevant information. The disentangled principle helps to reduce the redundancy of stored information and decompose the entanglement of the knowledge representations, which gives rise to the invariance of the produced knowledge vectors. EDK is model-agnostic, flexible, and tailored for fine-grained instance-level knowledge augmentation. Experiments and various ablation studies justify the effectiveness of it.

### v2

@@ 5.3.1. Impact of the Two Constraints
To enforce the parametric knowledge base to memorize the essential and disentangle knowledge patterns within limited model parameters, we propose two constraints to regularize the knowledge compression process. In this section, we investigate the effectiveness of the two proposed constraints. We conduct experiments with backbones on which FINED achieves the best performance, which is DeepFM for AD and DCN for eleme, respectively. We remove the regularizations of the essential and disentangled principles to see the impacts of them.
Table 4. Impacts of the two constraints.
Model AD Eleme
FINED 0.6378 0.1935 0.6563 0.0881
FINED w/o Disentangled 0.6356 0.1939 0.6433 0.0886
FINED w/o Essential 0.6347 0.1937 0.6520 0.0884
FINED w/o Both 0.6304 0.1940 0.6411 0.0888
@@
From Table 4, we can see that the proposed essential and disentangled constraints improve the quality of the memorized knowledge of the parametric knowledge base. Both the essential principle and the disentangled principle can solely contribute to the final performance, while the two complement each other and can collaborate for more qualified knowledge.
5.3.2. Impact of the Knowledge Extractor
@@
The patterns of the old data space are diverse and complex, being hard to be fully memorized due to the existence of spurious patterns and entanglement. To memorize the informative patterns within the data space, a knowledge extractor is designed to extract the essential and disentangled patterns within data points. In this section, we conduct experiments on the number of knowledge patterns per data point to study the impact of the knowledge extractor. Similarly, we select DeepFM as the backbone for AD dataset and DCN as the backbone for Eleme dataset. We test on the range of $\left[5,10,15,20,25\right]$, the results are listed in Figure 3.
@@
Figure 3. Performances of FINED w.r.t different number of knowledge patterns per sample.
@@
5.4. Case Study (RQ4)
5.4.1. Loss Curves of Different Objectives
@@
Figure 4. Learning curves of different losses on AD.
The learning curves of different losses on AD dataset is illustrated in Figure 4. The blue line displays the loss curve for the disentangled constraint, which is $\sum_{i\neq j}I(s_{i};s_{j})$. The red line and black line together make up the loss for the essential constraint, which is $-\alpha\sum_{j}I(\mathbf{s_{j}};y)+I(X;\mathbf{c})$. Concretely, the black line illustrates the curve for $-\sum_{j}I(\mathbf{s_{j}};y)$ which promotes the memorization for task relevant information, and the red line illustrates the curve for $I(X;\mathbf{c})$, which compresses the representation and helps to filter the noisy information. From the curves, we can see that all losses are learned well. While for the red one which corresponds to the vCLUB value, the MI upper bound estimation increases at the first epoch and decrease in the following training steps, since it takes some steps first to train the MI estimator well for correct estimation.
@@ 5.4.2. Visualizations of the Extracted Patterns
Then, we visualize the extracted patterns for more explicit analysis.
@@
Figure 5. Pattern scales. (Entry $>0.5$ is regarded as 1.)
@@
Firstly, we study the masks generated by the knowledge extractor. Note that the entries of the masks are continuous. Here we regard an entry with value larger than $0.5$ as an existing feature and calculate the number of the non-zero entries of the generated masks. The extracted pattern scale distribution is illustrated in Figure 5. From the results, we could find that the scales of extracted patterns vary, which validates the effectiveness of the knowledge extractor for extracting patterns of arbitrary scales.
In addition, we further study the representations of the raw data instance features and the extracted patterns. We use the T-SNE visualization to illustrate the distribution of the encodings of raw data instances and the extracted pattern vectors. The visualization results are displayed in Figure 6.Each point in the left denotes a representation of a raw data instance, and each point in the right represents an encoded pattern vector. From the results, we could see that there are apparently different manifolds existing in the extracted patterns representation space for both AD and Eleme datasets, while the raw data instances just appear in different clusterings. This validates the effectiveness of the proposed method on extracting the diverse and complex patterns from the old data.
@@ 6. Conclusion
In this paper, we design a parametric knowledge base FINED that compresses the massive old data into compact knowledge stored in parameters. Concretely, an instance-wise knowledge extractor is utilized to extract the patterns of arbitrary scales within data points and a knowledge encoder is utilized to memorize the patterns. The parametric knowledge base is regularized by the proposed essential and disentangled principles, which promote robust and generalized knowledge memorized by the parametric knowledge base without increasing the size of it. FINED is model-agnostic, flexible, and tailored for fine-grained instance-level knowledge augmentation.
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.
Appendix A Hyperparameter